OpenStudy (anonymous):

Function: Let f(x)= x^3 +4x^2 +Ax +2 and f(1)=15. Find the value of A.

jimthompson5910 (jim_thompson5910):

f(1)=15 means you plug in x = 1 and the output of the function is 15

jimthompson5910 (jim_thompson5910):

f(x)= x^3 +4x^2 +Ax +2 f(1)= (1)^3 +4*(1)^2 +A*(1) +2 ... replace every x with 1 15 = (1)^3 +4*(1)^2 +A*(1) +2 ... plug in f(1) = 15 now solve for A

OpenStudy (anonymous):

So it going to be 8? @jim_thompson5910

OpenStudy (anonymous):

So the value for A is 8 right?

jimthompson5910 (jim_thompson5910):

A = 8 is correct